A uniform solid sphere is rolling without slipping along a horizontal surface with a speed of 5.5 m/s when it starts up a ramp that makes an angle of 25.0° with the horizontal. What is the speed of the sphere after it has rolled 3.0 m up the ramp, measured along the surface of the ramp?
Answer #1

OG 3M 3m Elvidas no V; osoja Viz 5.5m/s 8= 250 d = 3m find speed of Sphere (VA)=? NE To fall kinefore energy 7 of solid spher7 21.195 = 0.7 V +(9.8113a sin25) 0.7 V = 21.145– 120437 V = 12-4819 TVE = 3053 m/s Ja Scanned with CamScarlehene er rolled